# MolGround: A Benchmark for Molecular Grounding

### v1

@@ 4.3 Structurization and Annotation
The structuring and annotation process is performed iteratively, allowing our grounding agent to collaborate with human chemical experts. Given the high cost of human intervention, we produced grounding results with two tiers of quality: 1) High-Quality Grounding Subset: This subset comprises 2% of the total dataset. For this portion, all substructures including rings, chains, atoms, and functional groups and their relationships have been manually validated with consensus from multiple human experts. 2) Coarse Grounding Subset: This subset accounts for the remaining 98% of the dataset, where all substructures and their relationships have been automatically generated without manual validation.
@@ 5.2 Evaluation of Pretrained Models
Table 3 compares the performance of LLMs and MLLMs across five molecular grounding tasks. Overall, most tasks remain challenging for all baseline models, with accuracy generally below 0.5. In particular, BNSM, SFA, and RSL prove to be the most difficult, with all models achieving accuracies below 0.337. By contrast, CNER and SRG exhibit relatively better performance. The highest F1-score for CNER is 0.633 and the highest accuracy for SRG is 0.803, which are achieved by GPT-4o. For the CNER task, models often over-extract functionality words as chemical substructure names (e.g., “amide functionality”) and struggle to correctly assign roles to extracted chemical names. For the SFA, models perform well on counting non-existent rings but struggle with monocyclic ring identification due to subtle structural differences. For more complex tasks (BNSM and SRG), performance remains low, even for the best model, GPT-4o, which achieves only 0.125 and 0.333 accuracy, respectively. In BNSM, models perform well for simple structures but struggle with complex mappings, often generating incorrect yet structurally similar names or SMILES strings. SRG results indicate difficulty in establishing substructures’ relationships, with models misinterpreting textual cues and neglecting structural connections. RSL task witnessed a particularly poor performance, with the best F1 (0.020) and IoU (0.174). MLLMs, despite access to structural images, do not surpass LLMs (except BNSM).
@@ 5.3 Evaluation of ICL and SFT
Table 4 examines how ICL and SFT impact the tasks. Overall, while both ICL and SFT improve results, the gains in RSL tasks remain limited. ICL significantly enhances tasks requiring chemical knowledge recall, particularly CNER and SFA. SFT provides the most substantial boost for grounding tasks. For instance, LLaMA 3.1-8B’s F1-score in singular RSL gains from 0.006 to 0.275, and in multiple RSL from 0.001 to 0.315. However, improvements are not consistent across models. ICL techniques sometimes degrade performance. For example, Few-shot learning and RAG lower GPT-4o’s accuracy on SRG, as additional examples introduce substructure relationships that distract the model.
@@ Appendix A Appendix
A.1 MolGround Data Split
We split the QAs into training, validation, and testing sets using a split ratio (80%, 10%, 10%) on each task, and ensure that there is no overlapped molecule between different sets.
A.2 Molecular Captioning Prompt Template
@@ A.3 Evaluation Metrics
A.4 Fine-grained Performance

### v2

@@ 4.3 Structurization and Annotation
The structuring and annotation process is performed iteratively, allowing our grounding agent to collaborate with human chemical experts. The 79k QA pairs listed in Table 1 are constructed based on the human annotation data. We split the QAs into training, validation, and testing sets using a split ratio (80%, 10%, 10%) on each task, and ensure that there is no overlapped molecule between different sets.
@@ 5.2 Evaluation of Pretrained Models
Table 3 compares the performance of LLMs and MLLMs across five molecular grounding tasks. Overall, most tasks remain challenging for all baseline models, with accuracy generally below 0.5. In particular, BNSM, SFA, and RSL prove to be the most difficult, with all models achieving accuracies below 0.337. By contrast, CNER and SRG exhibit relatively better performance. The highest F1-score for CNER is 0.633 and the highest accuracy for SRG is 0.803, which are achieved by GPT-4o. For the CNER task, models often over-extract functionality words as chemical substructure names (e.g., “amide functionality”) and struggle to correctly assign roles to extracted chemical names. For the SFA, models perform well on counting non-existent rings but struggle with monocyclic ring identification due to subtle structural differences. For example, the predicted probabilities of the benzene to be a donor or acceptor is Moreover, BioT5+ fails to follow instructions, and Mol-Instructions exhibits a strong bias in selecting question choices . While chemical name recognition (CNER) is relatively well-handled, with most models achieving an F1-score above 0.9, smaller models (e.g., LLaMA 3.1-8B, ChemLLM-7B) similar names (e.g., ether vs. ester) For more complex tasks (BNSM and SRG), performance remains low, even for the best model, GPT-4o, which achieves only 0.125 and 0.333 accuracy, respectively. In BNSM, models perform well for simple structures but struggle with complex mappings, often generating incorrect yet structurally similar names or SMILES strings. SRG results indicate difficulty in establishing substructures’ relationships, with models misinterpreting textual cues and neglecting structural connections. RSL task witnessed a particularly poor performance, with the best F1 (0.020) and IoU (0.174). In singular RSL, GPT-4o achieves the highest accuracy (0.755) but with low IoU (0.148), while LLaMA 3.1-8B prioritizes broader highlighting at the cost of accuracy. For multiple RSL, most models generate invalid outputs and show low substructure coverage (max 39.9%). MLLMs, despite access to structural images, do not surpass LLMs (except BNSM).
@@ 5.3 Evaluation of ICL and SFT
Table 4 examines how ICL and SFT impact the tasks. Overall, while both ICL and SFT improve results, the gains in RSL tasks remain limited. ICL significantly enhances tasks requiring chemical knowledge recall, particularly CNER and SFA. Few-shot learning boosts GPT-4o and LLaMA 3.1-8B accuracy to 99.4% and 96.0% on CNER, respectively. RAG dramatically improves SFA performance, with increases of 124% for GPT-4o and 489% for LLaMA 3.1-8B. Mol-Instructions sees an extreme gain of 56009%, though this may indicate prior poor performance rather than true capability improvement. SFT provides the most substantial boost for grounding tasks. For instance, LLaMA 3.1-8B’s F1-score in singular RSL gains from 0.006 to 0.275, and in multiple RSL from 0.001 to 0.315. Mol-Instructions follows a similar trend, demonstrating SFT’s effectiveness in refining molecular representations. However, improvements are not consistent across models. ICL techniques sometimes degrade performance. For example, Few-shot learning and RAG lower GPT-4o’s accuracy on SRG, as additional examples introduce substructure relationships that distract the model.
@@ Appendix A Appendix
A.1 Molecular Captioning Prompt Template
@@ A.2 Evaluation Metrics
A.3 Fine-grained Performance